# Accurate Checkerboard Corner Detection under Defoucs

### v1

@@ 5 Discussions
Conclusion In this paper, we proposed a novel feature extraction method for camera calibration, addressing the limitations of existing approaches, particularly for visible light camera cameras under defocus and distortion conditions. Our objective function, which accounts for blur nature of optical defocus, simplifies computation and reduces the risk of overfitting for symmetry based method proposed by Schöps et al.. Through experiments on real-world data from both infrared and visible light camera cameras, we demonstrated that our method significantly improves reprojection error compared to OpenCV and approach by Schöps et al.. Specifically, we achieved a mean and medium reprojection error reduction of and than Schöps et al. for visible light camera cameras with large distortion. For infrared case, we get a similar result with optimal refinement window size. In both types of cameras, we achieved significantly better results compared to OpenCV, reduce reprojection error to . We also discussed the impact of optimization window size on the algorithms and found that it is important for cameras with large distortion. In such cases, OpenCV is highly sensitive to the optimization window size, while our method is more robust.
We discussed two different causes of blurred images and explained why the method by Schöps et al. is not suitable for visible light cameras. For visible light cameras, we considered the Gaussian point spread function blur model for straight edges of checkerboard patterns and provided an analytical solution for discrete pixel values. This is significant for further calibration methods based on deconvolution, as well as precise localization in applications like QR code detection. Our calculations show that when defocus is small, there is a significant difference between point intensity and pixel intensity, which may lead to more accurate results in current single-image depth from defocus methods based on point intensity.
In conclusion, our method offers a significant improvement in feature extraction accuracy for camera calibration, especially in real-world scenarios where distortion, defocus, and challenging poses are present. These findings suggest that our approach is well-suited for practical applications in 3D vision, autonomous systems, and multi-sensor environments, where reliable and accurate calibration is critical.
Limitations and feature works For the visible light camera, we calibrated not only industrial cameras but also smartphone cameras. Although our method still achieved better results than OpenCV, the difference was not as significant as we expected. We observed noticeable purple fringing in the smartphone camera when capturing the checkerboard pattern, which may be caused by chromatic aberration. Therefore, there is room for algorithmic improvement for cheap digital cameras. Another issue is, although our experiments show that the method by Schöps et al. is not optimal for visible light cameras, it still achieved significantly better results than OpenCV with the star pattern they proposed. This is because, as explained by Datta et al., the assumptions of the OpenCV method only hold for non fronto-parallel input images, meaning at most two intersecting lines meet the assumptions, which limits its ability to effectively utilize the information from all lines. However, as reported by Schöps et al. increasing the number of intersection lines on the star pattern beyond does not further enhance accuracy. How to resolve aliasing at the center of star based pattern to achieve higher precision remains to be studied further. This will be a direction for our future work.

### v2

@@ 5 Discussions
Conclusion In this paper, we introduced a novel approach to camera calibration, focusing on both pattern and cost function design. Our method demonstrated significant improvements over previous techniques in terms of both accuracy and computational efficiency. Specifically, in a dataset with more than one thousand images provide by Schöps et al. for generic camera calibration, our cost function reduced the median reprojection error by compared to their method, while also decreasing computation time by more than . The proposed dynamic calibration pattern, based on rotating checkerboard sequences, further enhanced accuracy, achieving a reduction in reprojection error compared to traditional checkerboard-based methods with OpenCV feature refinement. Our method provides new insights for high-precision generic camera calibration and structured light design. We also discussed the impact of interpolation on existing symmetry-based cost functions through the analytical expression of the defocus model. We used a Gaussian kernel for the calculation, and give the analytical expression of pixel integral. Under typical parameters, our results show that when the Gaussian sigma is around just 1 pixel, the difference between pixel area integral and the pixel center can be ignored. Therefore, our cost function based on interpolation techniques can be effective in conventional camera calibration.
Limitations and feature works We tested our calibration pattern on the standard Brown-Conrady model, which inevitably introduces systematic errors. To apply it to a generic model, we need to further engineer and optimize the implementation of the proposed pattern to enhance its data acquisition speed. Additionally, we have experimentally validated that multiple checkerboard patterns can achieve higher calibration accuracy compared to the star-shaped pattern. However, we have not yet discussed the limits of pattern amounts. A quantitative idea is illustrated in the Fig. 11 bellow: we need to distinguish the two adjacent segments at the edge of the optimization window, ensuring they fall on different pixels. For a window with half-width of , at most segments can be resolved. However, we need to account for defocus and interpolation effects, which make the precise computation of results difficult, and this requires further research.
There is one issue that should be mentioned again, although we used a screen to display our calibration pattern, it does not rely on the display device in the same way as other dynamic markers. In fact, if we change the bounding shape of our pattern from square to circular, we can manufacture it mechanically and drive it with motors. This would make it applicable in situations where a display cannot be used, and this will be the focus of our next experimental direction.